Silver tarnishes in the presence of hydrogen sulfide (which smells like rotten eggs) and oxygen because of the reaction
4 Ag + 2 H2S + O2  2 Ag2S + 2 H2O
How many grams of silver sulfide can be formed from 1.1 g Ag?
1
Expert's answer
2020-04-09T13:11:29-0400

4Ag + 2H2S + O2 = 2Ag2S + 2H2O


"n(Ag)=\frac{m}{M}=frac{1.1}{107.8682}=0.0102"


"\frac{0.0102}{4}=\frac{x}{2}"


x=0.005


"m(Ag2S)=n\times M=0.005\times247.8024=1.26"